# Adversarial Attacks of Vision Tasks in the Past 10 Years: A Survey

### v1

@@ 5.1.4. Generalization Encouraging
Generalization can be improved by generative model training or aggregation. The training of generative models is essentially a process of aggregating and generalizing over a dataset. The difference between the two lies in their targets: the former targets the model itself, while the latter targets the perturbations.
Generative models can be either pre-trained or not. The former leverages the inherent generalization ability of the pre-trained model to enhance the naturalness of AEs and uses gradient aggregation to improve transferability. The latter directly trains the generative model to learn adversarial semantics for generating adversarial perturbations. AdvDiffVLM (Guo et al., 2024), using a vision-language model (VLM) as the victim, employs a pre-trained diffusion model to enhance sample naturalness, while gradient aggregation embeds adversarial semantics to improve transferability. GAP (Poursaeed et al., 2018), AdvEncoder (Zhou et al., 2023b), and AdvCLIP (Zhou et al., 2023a) respectively train autoencoders, generators, and GANs to generate universal perturbations, achieving cross-image capabilities via generative models. Although generative-model-based methods may add an extra training stage, they often offer a speed advantage during perturbation generation.
Aggregation-based methods can be categorized into: 1) dataset aggregation (Moosavi-Dezfooli et al., 2017), 2) model aggregation (Liu et al., 2016; Xiong et al., 2022; Huang et al., 2023a), and 3) transformation aggregation (Xie et al., 2019b; Dong et al., 2019; Lin et al., 2019; Xie et al., 2017b; Yan et al., 2022; Long et al., 2022; Wang et al., 2021).
@@
Model aggregation. Based on the assumption, if a sample can be misclassified by multiple models, it may also be misclassified by others (Liu et al., 2016), model aggregation improves transferability by combining the loss (Dong et al., 2018), confidence scores (Liu et al., 2016), or logits (Dong et al., 2018; Xie et al., 2019b; Chen et al., 2023c; Xiong et al., 2022) from multiple surrogate models. Inspired by stochastic depth (Huang et al., 2016), T-SEA (Huang et al., 2023a) does not aggregate across different surrogate models but instead implicitly ensembles the variants of the single model by randomly dropping some sub-layers.
Transformation aggregation. Similar to dataset aggregation, transformation aggregation (also known as data augmentation) improves transferability by stacking multiple inputs. The difference is that transformation aggregation stacks different transformed versions of the same sample, such as through translation (Dong et al., 2019), scaling (Lin et al., 2019; Xie et al., 2019b), padding (Xie et al., 2019b), etc. Additionally, DAG (Xie et al., 2017b) achieves an effect similar to translation transformation aggregation by combining gradients from multiple proposal boxes on the same target. ILA-DA (Yan et al., 2022) sets learnable transformation selection parameters, which adaptively sample transformation functions like translation, shearing, and rotation during iterations and apply them to the sample for data augmentation. SSA (Long et al., 2022) simulates model aggregation by averaging gradients from multiple random DCT and IDCT (Ahmed et al., 1974) samples (random DCT refers to adding random noise to the sample after DCT). Admix (Wang et al., 2021) attempts to combine transformation aggregation and dataset aggregation: on the basis of scaled transformation aggregation, it mixes the samples to be aggregated with a small portion of samples from other classes before computing the perturbation. Transform aggregation is the most notable among the three aggregation methods.
@@ 5.2. Improving Physical Robustness
Physical robustness helps AEs retain effectiveness across device environments like smartphones, cameras, or printed formats (see section 2.3). Creating physically robust AEs is harder than standard ones, as they must consider the affects on various distances, angles, lighting, and camera constraints. Two main approaches are used: 1) designing specific objective functions (Sharif et al., 2016; Thys et al., 2019; Komkov and Petiushko, 2021), and 2) physical transformation aggregation (Sharif et al., 2016; Thys et al., 2019; Komkov and Petiushko, 2021; Athalye et al., 2018; Brown et al., 2017; Eykholt et al., 2018; Chen et al., 2019; Song et al., 2018; Liu et al., 2020; Huang et al., 2024; Zheng et al., 2023) (e.g., varying position, distance, angle, lighting, background, etc.). For the objective function, TV loss (Rudin et al., 1992) can be used to smooth the perturbations or constrain pixel values within printable ranges. In transformation aggregation, a transformation distribution, aka Expectation Over Transformation (EOT) (Athalye et al., 2018), is often constructed. This distribution simulates various real-world conditions that an image might encounter. AEs are then randomly transformed, and the gradients from these transformed samples are aggregated to make samples robust to physical conditions. Additionally, BIM (Kurakin et al., 2018) finds that fine-grained perturbations are easily disrupted during capture, making the single-step FGSM (Goodfellow et al., 2014) more robust than iterative BIM. AGN (Sharif et al., 2019) trains GANs with the transformed to generate physically robust eyeglasses.
@@ 5.3. Improving Stealthiness
Stealthiness makes AEs invisible to humans or detectors. Methods to improve stealthiness fall into four types: 1) limiting perturbation size (Kurakin et al., 2018; Madry, 2017; Moosavi-Dezfooli et al., 2017; Gao et al., 2020b, a; Carlini and Wagner, 2017; Chen et al., 2017; Kos et al., 2018), 2) restricting perturbation areas (Papernot et al., 2016b; Su et al., 2019; Narodytska and Kasiviswanathan, 2016), 3) disguising perturbations (Brown et al., 2017; Eykholt et al., 2018; Song et al., 2018; Komkov and Petiushko, 2021; Liu et al., 2020; Li et al., 2019; Zheng et al., 2023; Jia et al., 2020), and 4) enhancing naturalness (Chen et al., 2023a; Guo et al., 2024; Liu et al., 2023c; Laidlaw et al., 2020; Zhao et al., 2020; Laidlaw and Feizi, 2019; Shamsabadi et al., 2021, 2020). The most common method is perturbation size limitation, where iterative attacks usually use clipping (Kurakin et al., 2018; Madry, 2017; Moosavi-Dezfooli et al., 2017) or projection (Gao et al., 2020b, a) to control the intensity, and optimization-based methods often minimize the perturbation norm (Carlini and Wagner, 2017; Chen et al., 2017; Kos et al., 2018). Unlike global perturbations, some search-based methods (Papernot et al., 2016b; Su et al., 2019; Narodytska and Kasiviswanathan, 2016) focus on finding a few vulnerable pixel locations for the attack, improving stealth by limiting the affected area.
@@
When perturbations can’t be hidden (e.g., universal patched or physically robust ones), disguising them as tie-dye (Brown et al., 2017), graffiti (Eykholt et al., 2018), stickers (Eykholt et al., 2018; Song et al., 2018; Komkov and Petiushko, 2021; Liu et al., 2020; Li et al., 2019; Zheng et al., 2023), or watermarks (Jia et al., 2020) is common, creating the illusion of art or pranks. While the above methods can still be detected sometimes, improving naturalness makes perturbations more invisible. These methods (Chen et al., 2023a; Guo et al., 2024; Liu et al., 2023c; Laidlaw et al., 2020; Zhao et al., 2020; Laidlaw and Feizi, 2019; Shamsabadi et al., 2021, 2020) do not limit the perturbation size but aim to make them imperceptible to human perception. As a result, the perturbations may take on some form based on the original sample, rather than resembling random noise (e.g., altering the color of a dog’s eyes or the texture of a spider’s back (Chen et al., 2023a)).
@@
The last method above generates one type of UAEs (Brown et al., 2018) (Unrestricted AEs). When generating RAEs (Restricted AEs), attackers are constrained, such as making limited modifications or following certain attack paradigms. In contrast, UAEs can be generated using any method, including applying larger perturbations, spatial transformations (Brown et al., 2018), color adjustments (Laidlaw and Feizi, 2019; Shamsabadi et al., 2021, 2020), limiting perceptual distance (Laidlaw et al., 2020; Zhao et al., 2020), or processing by generative models (Chen et al., 2023a; Guo et al., 2024). UAEs tend to test model robustness and RAEs are more realistic, while the fourth method mentioned here balances both.
5.4. Improving Speed
@@
Improving the speed of AEs generation enables quicker attacks, achieved through three main methods: 1) single-step approaches (Goodfellow et al., 2014; Tramèr et al., 2017b, a), 2) fast-converging algorithms (Zhang et al., 2020; Pintor et al., 2021; Croce and Hein, 2020a; Wu et al., 2020), and 3) generative models (Baluja and Fischer, 2017; Poursaeed et al., 2018; Xiao et al., 2018; Zhou et al., 2023b, a; Chen et al., 2023a; Guo et al., 2024). Single-step methods, though fast, may have limited effectiveness. BP (Zhang et al., 2020) and FMN (Pintor et al., 2021) accelerate convergence by reducing search oscillation via parameter adjustments and cosine annealing, respectively, while FAB (Croce and Hein, 2020a) and SWFA (Wu et al., 2020) improve efficiency with precise projection techniques. FMN also optimizes input space exploration with adaptive box constraints, and SWFA accelerates calculations through sparsity (limiting the perturbation range) and gradient normalization (better direction). Generative models, though requiring a training phase, can generate examples with a single forward pass (Baluja and Fischer, 2017; Poursaeed et al., 2018; Xiao et al., 2018). Diffusion models, despite needing iterative denoising to inject adversarial semantics, still outperform popular iterative methods in speed (Chen et al., 2023a; Guo et al., 2024).
6. Adversarial Attacks beyond Classification
@@
Due to the pervasive nature of AEs, attack techniques cover various tasks, making it hard to list all scenarios. As Qi et al. (Qi et al., 2024) noted, the rise of LLMs has shifted adversarial attacks from a classification-centric focus to a broader approach encompassing all LLM applications. Traditional tasks include classification-centric activities such as recognition (Szegedy, 2013; Dong et al., 2018; Chen et al., 2023a), detection (Xie et al., 2017b; Huang et al., 2023a; Thys et al., 2019), segmentation (Cisse et al., 2017; Li et al., 2020a; Poursaeed et al., 2018), and reinforcement learning (Huang et al., 2017; Lin et al., 2017), as well as generative tasks like image generation (Pasquini et al., 2019), translation (Kos et al., 2018; Tabacof et al., 2016; Gondim-Ribeiro et al., 2018), and super-resolution (Yin et al., 2018). In cases of infringement, such as style transfer (Johnson et al., 2016) and face swapping (Bode et al., 2021), these attacks help protect artists’ copyrights (Shan et al., 2023) and degrade swap quality (Dong et al., 2023b). Notably, Jia et al. (Jia et al., 2020) have used watermarking to create visually meaningful perturbations (as opposed to meaningless noise), which have potential applications in data copyright protection. Recently, research has gradually shifted toward multimodal tasks with LVLMs like image captioning (Chen et al., 2023c; Guo et al., 2024; Dong et al., 2023a; Zhao et al., 2024; Wang et al., 2023b; Schlarmann and Hein, 2023), visual question answering (VQA) (Luo et al., 2024a; Cui et al., 2024; Zhao et al., 2024; Wang et al., 2023b; Schlarmann and Hein, 2023; Tu et al., 2023; Wang et al., 2024b), and vision-language retrieval (Zhou et al., 2023a; Zheng et al., 2024; Wang et al., 2024a), highlighting new directions for adversarial attacks.

### v2

@@ 5.1.4. Generalization Encouraging
Improving generalization helps to enhance transferability, which can be achieved by generative model training (Poursaeed et al., 2018; Zhou et al., 2023b, a) or aggregation (Moosavi-Dezfooli et al., 2017; Liu et al., 2016; Xie et al., 2019b). The training of generative models is essentially a process of aggregating and generalizing over a dataset. The difference between the two lies in their targets: the former targets the model itself, while the latter targets the perturbations.
Generative models can be either pre-trained or not. The former leverages the inherent generalization ability of the pre-trained model to enhance the naturalness of AEs and uses gradient aggregation to improve transferability. The latter directly trains the generative model to learn adversarial semantics for generating adversarial perturbations. AdvDiffVLM (Guo et al., 2024), using a vision-language model (VLM) as the victim, employs a pre-trained diffusion model to enhance sample naturalness, while gradient aggregation embeds adversarial semantics to improve transferability. GAP (Poursaeed et al., 2018), AdvEncoder (Zhou et al., 2023b), and AdvCLIP (Zhou et al., 2023a) respectively train autoencoders, generators, and GANs to generate universal perturbations, achieving cross-image capabilities via generative models. Although generative-model-based methods may add an extra training stage, they often offer a efficiency advantage during perturbation generation.
Aggregation-based methods can be categorized into: 1) dataset aggregation (Moosavi-Dezfooli et al., 2017), 2) model aggregation (Liu et al., 2016; Xiong et al., 2022; Huang et al., 2023a), and 3) transformation aggregation (Xie et al., 2019b; Dong et al., 2019; Lin et al., 2020; Xie et al., 2017b; Yan et al., 2022; Long et al., 2022; Wang et al., 2021c).
@@
Model aggregation. Based on the assumption, if a sample can be misclassified by multiple models, it may also be misclassified by others (Liu et al., 2016), model aggregation improves transferability by combining the loss (Dong et al., 2018), confidence scores (Liu et al., 2016), or logits (Dong et al., 2018; Xie et al., 2019b; Chen et al., 2023d; Xiong et al., 2022) from multiple surrogate models. Inspired by stochastic depth (Huang et al., 2016), T-SEA (Huang et al., 2023a) does not aggregate across different surrogate models but instead implicitly ensembles the variants of the single model by randomly dropping some sub-layers.
Transformation aggregation. Similar to dataset aggregation, transformation aggregation (also known as data augmentation) improves transferability by stacking multiple inputs. The difference is that transformation aggregation stacks different transformed versions of the same sample, such as through translation (Dong et al., 2019), scaling (Lin et al., 2020; Xie et al., 2019b), padding (Xie et al., 2019b), etc. Additionally, DAG (Xie et al., 2017b) achieves an effect similar to translation transformation aggregation by combining gradients from multiple proposal boxes on the same target. ILA-DA (Yan et al., 2022) sets learnable transformation selection parameters, which adaptively sample transformation functions like translation, shearing, and rotation during iterations and apply them to the sample for data augmentation. SSA (Long et al., 2022) simulates model aggregation by averaging gradients from multiple random DCT and IDCT (Ahmed et al., 1974) samples (random DCT refers to adding random noise to the sample after DCT). Admix (Wang et al., 2021c) attempts to combine transformation aggregation and dataset aggregation: on the basis of scaled transformation aggregation, it mixes the samples to be aggregated with a small portion of samples from other classes before computing the perturbation. Transform aggregation is the most notable among the three aggregation methods.
@@ 5.2. Improving Physical Robustness
Physical robustness helps AEs retain effectiveness across environments. Creating physically robust AEs is harder than standard ones, as they must consider the affects on various distances, angles, lighting, and device constraints (section 2.3). Two main approaches are used: 1) designing specific objective functions (Sharif et al., 2016; Thys et al., 2019; Komkov and Petiushko, 2021), and 2) physical transformation aggregation (Sharif et al., 2016; Thys et al., 2019; Komkov and Petiushko, 2021; Athalye et al., 2018; Brown et al., 2017; Eykholt et al., 2018; Chen et al., 2019; Song et al., 2018; Liu et al., 2020; Huang et al., 2024; Zheng et al., 2023). For the objective function, TV loss (Rudin et al., 1992) can be used to smooth the perturbations or constrain pixel values within printable ranges. In transformation aggregation, a transformation distribution, aka Expectation Over Transformation (EOT) (Athalye et al., 2018), is often constructed. This distribution simulates various real-world conditions that an image might encounter. AEs are then randomly transformed, and the gradients from these transformed samples are aggregated to make samples robust to physical conditions. Additionally, BIM (Kurakin et al., 2018) finds that fine-grained perturbations are easily disrupted during capture, making the single-step FGSM (Goodfellow et al., 2014) more robust than iterative BIM. AGN (Sharif et al., 2019) trains GANs with the transformed to generate physically robust eyeglasses.
@@ 5.3. Improving Stealthiness
Methods to improve stealthiness fall into four types: 1) limiting perturbation size (Kurakin et al., 2018; Madry, 2017; Moosavi-Dezfooli et al., 2017; Gao et al., 2020b, a; Carlini and Wagner, 2017; Chen et al., 2017; Kos et al., 2018), 2) restricting perturbation areas (Papernot et al., 2016c; Su et al., 2019; Narodytska and Kasiviswanathan, 2016), 3) disguising perturbations (Brown et al., 2017; Eykholt et al., 2018; Song et al., 2018; Komkov and Petiushko, 2021; Liu et al., 2020; Li et al., 2019; Zheng et al., 2023; Jia et al., 2020), and 4) enhancing naturalness (Chen et al., 2023b; Guo et al., 2024; Liu et al., 2023c; Laidlaw et al., 2020; Zhao et al., 2020; Laidlaw and Feizi, 2019; Shamsabadi et al., 2021, 2020). The most common method is perturbation size limitation, where iterative attacks use clipping (Kurakin et al., 2018; Madry, 2017; Moosavi-Dezfooli et al., 2017) or projection (Gao et al., 2020b, a), and optimization-based ones minimize perturbation norms (Carlini and Wagner, 2017; Chen et al., 2017; Kos et al., 2018). Constraints can be enforced via norms or perceptual metrics like SSIM and LPIPS. However, norm-constrained perturbations often appear as irregular, noise-like patterns, making them easily perceptible (Jordan et al., 2019; Wang et al., 2021a; Hameed and Gyorgy, 2021; Karli et al., 2022; Clements and Lao, 2022). In contrast, perceptual distance constraints enable coherent, semantically meaningful perturbations, improving AE fidelity (Wang et al., 2021a; Hameed and Gyorgy, 2021; Karli et al., 2022; Clements and Lao, 2022). Additionally, some search-based methods (Papernot et al., 2016c; Su et al., 2019; Narodytska and Kasiviswanathan, 2016) improve stealthiness by identifying sparse vulnerable pixels, minimizing the attack’s spatial footprint.
@@
When perturbations can’t be hidden (e.g., universal patched or physically robust ones), disguising them as tie-dye (Brown et al., 2017), graffiti (Eykholt et al., 2018), stickers (Eykholt et al., 2018; Song et al., 2018; Komkov and Petiushko, 2021; Liu et al., 2020; Li et al., 2019; Zheng et al., 2023), or watermarks (Jia et al., 2020) is common, creating the illusion of art or pranks. While the above methods can still be detected sometimes, improving naturalness makes perturbations more invisible. These methods (Chen et al., 2023b; Guo et al., 2024; Liu et al., 2023c; Laidlaw et al., 2020; Zhao et al., 2020; Laidlaw and Feizi, 2019; Shamsabadi et al., 2021, 2020) do not limit the perturbation size but aim to make them imperceptible to human perception. As a result, the perturbations may take on some form based on the original sample, rather than resembling random noise (e.g., altering the color of a dog’s eyes or the texture of a spider’s back (Chen et al., 2023b)).
@@
Improving naturalness generates one type of UAEs (Brown et al., 2018) (Unrestricted AEs). When generating RAEs (Restricted AEs), attackers are constrained, such as making limited modifications or following certain attack paradigms. In contrast, UAEs can be generated using any method, including applying larger perturbations, spatial transformations (Brown et al., 2018), color adjustments (Laidlaw and Feizi, 2019; Shamsabadi et al., 2021, 2020), limiting perceptual distance (Laidlaw et al., 2020; Zhao et al., 2020), or processing by generative models (Chen et al., 2023b; Guo et al., 2024). UAEs tend to evaluate model robustness and RAEs are more realistic, while the AEs from generative models balances both.
5.4. Increasing Efficiency
@@
Increasing the efficiency of AEs generation enables quicker attacks, achieved through three main methods: 1) single-step approaches (Goodfellow et al., 2014; Tramèr et al., 2017b, a), 2) fast-converging algorithms (Zhang et al., 2020; Pintor et al., 2021; Croce and Hein, 2020; Wu et al., 2020), and 3) generative models (Baluja and Fischer, 2017; Poursaeed et al., 2018; Xiao et al., 2018; Zhou et al., 2023b, a; Chen et al., 2023b; Guo et al., 2024). Single-step methods, though fast, may have limited effectiveness. BP (Zhang et al., 2020) and FMN (Pintor et al., 2021) accelerate convergence by reducing search oscillation via parameter adjustments and cosine annealing, respectively, while FAB (Croce and Hein, 2020) and SWFA (Wu et al., 2020) improve efficiency with precise projection techniques. FMN also optimizes input space exploration with adaptive box constraints, and SWFA accelerates calculations through sparsity (limiting the perturbation range) and gradient normalization (better direction). Generative models, though requiring a training phase, can generate examples with a single forward pass (Baluja and Fischer, 2017; Poursaeed et al., 2018; Xiao et al., 2018). Diffusion models, despite needing iterative denoising to inject adversarial semantics, still outperform popular iterative methods in efficiency (Chen et al., 2023b; Guo et al., 2024).
5.5. Improving Interpretability of Results
In high-risk domains such as medical diagnosis and financial risk control, merely manipulating model outputs is insufficient—successfully deceiving users requires additionally misleading interpretation mechanisms (e.g., Grad-CAM heatmaps (Selvaraju et al., 2017)). Such attacks fall into two categories: output-interpretation mismatch attacks and match attacks. Mismatch attacks create conflicts between outputs and interpretations to induce user distrust or misjudgment. For instance, a chest X-ray showing pneumonia in the left lung might be correctly classified as “pneumonia,” yet Grad-CAM highlights the right lung, potentially misleading clinicians. Match attacks induce model errors while maintaining output-interpretation consistency. For example, a pneumonia-free X-ray might be misclassified as “pneumonia” with supporting Grad-CAM evidence. Vadillo et al. (Vadillo et al., 2025) categorized eight adversarial attack scenarios based on varying degrees of output and interpretation errors, systematically analyzing model interpretation-aware attacks.
6. Adversarial Examples in Diverse Applications and Modalities.
every leaf node/.style=
if n children=0#1
every tree node/.style=
if n children=0minimum width=1em#1
{forest}
nonleaf/.style=font=,
for tree=every leaf node=my leaf, font=,
every tree node=my node, font=, l sep-=4.5pt, l-=1.pt,
anchor=west,
inner sep=2pt,
l sep=10pt, s sep=3pt, fit=tight,
grow’=east,
edge=ultra thin,
parent anchor=east,
child anchor=west,
if n children=0nonleaf,
edge path=
[…]
if=isodd(n_children())
for children=
if=equal(n,(n_children("!u")+1)/2)calign with current
Applications and Modalities
(section 6), draw=brightlavender, fill=brightlavender!15, text width=3.2cm, text=black
Applications of AEs, draw=brightlavender, fill=brightlavender!15, text width=3cm, text=black
Robustness Evaluation, draw=brightlavender, fill=brightlavender!15, text width=4cm, text=black
[(Carlini and Wagner, 2017; Carlini et al., 2024; Szegedy et al., 2014; Pei et al., 2017) , color=brightlavender, fill=brightlavender!15, text width=1.9cm, text=black]
Designing Robust Systems, draw=brightlavender, fill=brightlavender!15, text width=4cm, text=black
[(Madry, 2017; Chen et al., 2023b) , color=brightlavender, fill=brightlavender!15, text width=1.9cm, text=black]
Interpretable Machine Learning, draw=brightlavender, fill=brightlavender!15, text width=4cm, text=black
[(Fawzi et al., 2017; Su et al., 2019; Dong et al., 2017) , color=brightlavender, fill=brightlavender!15, text width=1.9cm, text=black]
Copyright Protection, draw=brightlavender, fill=brightlavender!15, text width=4cm, text=black
[(Shan et al., 2023; Dong et al., 2023b) , color=brightlavender, fill=brightlavender!15, text width=1.9cm, text=black]
Modal-Specific AEs, draw=brightlavender, fill=brightlavender!15, text width=3cm, text=black
Vision, draw=brightlavender, fill=brightlavender!15, text width=4cm, text=black
[(Cisse et al., 2017; Huang et al., 2017; Pasquini et al., 2019; Kos et al., 2018; Yin et al., 2018)… , color=brightlavender, fill=brightlavender!15, text width=1.9cm, text=black]
Language, draw=brightlavender, fill=brightlavender!15, text width=4cm, text=black
[(Zou et al., 2023; Wei et al., 2024a) , color=brightlavender, fill=brightlavender!15, text width=1.9cm, text=black]
Speech, draw=brightlavender, fill=brightlavender!15, text width=4cm, text=black
[(Cisse et al., 2017; Bahramali et al., 2021) , color=brightlavender, fill=brightlavender!15, text width=1.9cm, text=black]
Multimodality, draw=brightlavender, fill=brightlavender!15, text width=4cm, text=black
[(Wang et al., 2024a; Lu et al., 2024; Li et al., 2024b; Gong et al., 2023b; Shayegani et al., 2023)… , color=brightlavender, fill=brightlavender!15, text width=1.9cm, text=black]
[…]
Fig. 7 illustrates variegated applications and modality extensions of AEs beyond image classification. In application scenarios, (Carlini and Wagner, 2017; Carlini et al., 2024; Szegedy et al., 2014; Pei et al., 2017) use AEs to assess the lower bounds of model performance under attacks achieving robustness evaluation of models, while (Madry, 2017; Chen et al., 2023b) design more robust systems through adversarial training. Additionally, AEs have been found to help describe the shape of decision boundaries (Fawzi et al., 2017), or provide geometric insights into the model’s input space (Su et al., 2019; Moosavi-Dezfooli et al., 2017). In cases of infringement, such as style transfer (Johnson et al., 2016) and face swapping (Bode et al., 2021), AEs help degrade swap quality (Dong et al., 2023b) and protect artists’ copyrights (Shan et al., 2023) (though (Hönig et al., 2024) find this may not reliably work). Notably, Jia et al. (Jia et al., 2020) have used watermarking to create visually meaningful perturbations (as opposed to meaningless noise), which have potential applications in data copyright protection.
@@ 5.5. Improving Interpretability of Results
For modality extensions, as Qi et al. (Qi et al., 2024) noted, the rise of LLMs has shifted adversarial attacks from a classification-centric focus to a broader facet encompassing all LLM applications. Traditional vision tasks include classification-centric activities such as recognition (Szegedy et al., 2014; Dong et al., 2018; Chen et al., 2023b), detection (Xie et al., 2017b; Huang et al., 2023a; Thys et al., 2019), segmentation (Cisse et al., 2017; Li et al., 2020b; Poursaeed et al., 2018), and reinforcement learning (Huang et al., 2017; Lin et al., 2017), as well as generative tasks like image generation (Pasquini et al., 2019), translation (Kos et al., 2018; Tabacof et al., 2016; Gondim-Ribeiro et al., 2018), and super-resolution (Yin et al., 2018). Beyond visual models, language (Zou et al., 2023; Wei et al., 2024a) and speech (Cisse et al., 2017; Bahramali et al., 2021) models are also susceptible to AEs. Recently, research has gradually shifted toward multimodal tasks with LVLMs like image captioning (Chen et al., 2023d; Guo et al., 2024; Dong et al., 2023a; Zhao et al., 2024; Wang et al., 2023b; Schlarmann and Hein, 2023), visual question answering (VQA) (Luo et al., 2024a; Cui et al., 2024; Zhao et al., 2024; Wang et al., 2023b; Schlarmann and Hein, 2023; Tu et al., 2023; Wang et al., 2024d), and vision-language retrieval (Zhou et al., 2023a; Zheng et al., 2024; Wang et al., 2024a), highlighting new directions for adversarial attacks. LVLM attack is a pivotal extension of the traditional attack.